OpenStudy (anonymous):

2x2 + 9x + 4 = 0 solve by using the quadratic formula.

OpenStudy (anonymous):

\[ax^2+bx+c=0\implies x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]Work it out for yourself.

OpenStudy (anonymous):

just substitute a,b and c into the eqn provided by bendt... here a=2 b=9 c=4

OpenStudy (radar):

and verify the results by factoring.

OpenStudy (anonymous):

Aww thanks you're awesome:)

OpenStudy (anonymous):

That was another method of finding the roots... Sorry, i deleted it thinking it might confuse you...

OpenStudy (anonymous):

It's okay you still tried to help so that's all that matters:)